OpenStudy (anonymous):

Confusssed, help please. using Substitution method 3x=4y=13 x=y-5

jimthompson5910 (jim_thompson5910):

first equation has too many equal signs

OpenStudy (precal):

3x+4y=13 perhaps 3x-4y=13 perhaps

jimthompson5910 (jim_thompson5910):

or maybe 3x=4y+13?

OpenStudy (precal):

yes, another possibility

OpenStudy (anonymous):

oops soorry its suppose to be 3x+4y=13

jimthompson5910 (jim_thompson5910):

Start with 3x+4y=13 and replace 'x' with y-5 3x+4y=13 3(y-5)+4y=13 From here, solve for y 3(y-5)+4y=13 3y-15+4y=13 7y-15=13 7y=13+15 7y=28 y=28/7 y=4 Then use this value of y to find x. x=y-5 x=4-5 x=-1 So x = -1 and y = 4